SPECTROSCOPY OF THE GROUND, FIRST AND SECOND EXCITED TORSIONAL STATES OF AC-ETALDEHYDE FROM 0.05 TO 1.6 THz.

\underline{VADIM V. ILYUSHIN}^{a}, IVAN SMIRNOV, EUGENE A. ALEKSEEV, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; LAURENT MARGULÈS^{b}, RO-MAN A. MOTIYENKO, Laboratoire de Physique des Lasers, Atomes et Molècules, UMR 8523 CNRS-Universitè Lille 1, Bâtiment P5, F-59655 Villeneuve d’Ascq Cedex, France; BRIAN DROUIN, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109-8099, USA.

A new global study of the acetaldehyde (CH_{3}CHO) spectrum is reported. The new measurements cover the frequency range from 49 GHz to 1.6 THz and have been carried out using three different spectrometers in IRA NASU (Ukraine), PhLAM Lille (France), and JPL (USA). The rotational transitions belonging to the three lowest torsional states, as well as previously published data on the FIR torsional bands, of the molecule have been analyzed using the rho-axis-method. The dataset consisting of more than 19700 line frequencies and including rotational transitions with J up to 66 and K_{a} up to 22 was fit using a model consisting of 117 parameters and weighted root-mean-square deviation of 0.71 has been achieved. Details of this new study and problems encountered in analysis of the second torsional state will be discussed.

aPart of this work was done within the Ukrainian-French CNRS-PICS 6051 project.

bThis work was supported by the CNES and the Action sur Projets de l’INSU, PCMI.